# revtex4-2Repair the float

### v1

@@ I Introduction
Recent advancements in quantum computing have invigorated research in combinatorial optimization, a field ripe with prospects for practical applications. The enthusiasm is evident in the intensive theoretical investigations dedicated to evolving quantum methodologies. Despite theoretical insights driving new quantum algorithms, particularly through worst-case asymptotic runtime analysis [1, 2, 3], demonstrating quantum advantages in real-world applications presents ongoing challenges. Combinatorial optimization problems serve as a focal point for this research [4, 5], offering diverse applications. However, they face stiff competition from powerful classical solvers such as Gurobi, CPLEX, and CP-SAT, which routinely solve problems with thousands of variables to provable optimality. A promising pursuit in this context involves developing benchmarking strategies [9, 10, 5, 11]. By integrating classical emulation techniques, quantitative complexity analyses, and algorithm-specific simulations, researchers aim to juxtapose quantum algorithms against these classical titans. This approach is vital for evaluating performance on problem instances that exceed the 50-qubit limit inherent in state vector-based quantum simulations.
@@ I-A Assumptions on the Quantum Processing Unit
The resource estimation setup for applying the QTG method to QKP and MDKP relies on several fault-tolerant assumptions for the quantum processing unit (QPU). It assumes that all qubits and gates are noiseless, logical components. The universal gate set includes all single-qubit gates, singly-controlled single-qubit rotations, and Toffoli gates, all of which have equal implementation cost. Disjoint gates can be executed in parallel, counting as a single cycle on the QPU. Singly-controlled single-qubit rotations and Toffoli gates can be applied to arbitrary tuples of quubits. Additionally, all multi-controlled gates have access to the same ancilla qubits and implementing a control on 00 is just as costly as on $1$. Finally, to calculate a runtime from the cycle counts, we assume a quantum cycle time limit of $1ns$ given by [14].
@@ I-B Related Work
The 00-$1$ Quadratic Knapsack Problem was first introduced by Gallo et al. [15] and has a straightforward quadratic programming formulation (see Section II). Its practical solvability often depends on instance properties such as density [16, 17]. This has drawn significant interest from the optimization community, leading to recent exact algorithms like those in [18, 19]. In contrast, the 00-$1$ Multidimensional Knapsack Problem tends to be more tractable for integer programming solvers such as Gurobi [20]. Recent approaches, such as those in [21], have proven particularly effective for solving large instances [17].
Apart from QTG, many other quantum algorithms and frameworks for solving combinatorial optimization problems have been proposed and studied, such as Quantum Annealing [22], Nested Quantum Search (NQS) [23], Quantum Approximate Optimization Algorithm (QAOA) [24], and Quantum Branch-and-Bound (QBnB) [3]. While NQS and QBnB theoretically promise an advantage over classical solvers, Wilkening et al. [5] showed they are significantly outperformed by both the QTG search method and classical solvers on large real-world benchmarks when applied to the 0-1 Knapsack Problem.
@@ Methods
The QTG $\mathcal{G}$ for the QKP follows the same design criteria as its counterpart for the ordinary 00-$1$ Knapsack Problem [5]: It is comprised of $n$ layer unitaries $U_{m}=U_{m}^{3}U_{m}^{2}U_{m}^{1}$ that correspond to including/excluding each item in superposition while keeping track of the remaining capacity and the total profit arising from both the diagonal profits $p_{m}$ and the off-diagonal profits $p_{mm^{\prime}}$. Throughout this process, feasibility is ensured by controlling the branching at each item $m$ on the comparison “$w_{m}\leq c_{\bm{x}}$”. That is, we create a superposition of item $m$ being excluded/included if and only if its weight does not surpass the remaining capacity stored in the second register $\mathcal{H}_{2}$. Correspondingly, if the latter expression is not valid the $m$-th qubit stays in the |$\ket{0}$-state, indicating the item’s exclusion. The corresponding unitary is denoted $U_{m}^{1}=C_{\geq w_{m}}^{2}(R_{y,b}^{x_{m}})$, where the $R_{y}$ gate is responsible for creating a biased superposition of item exclusion (|$\ket{0}$-state) and inclusion (|$\ket{1}$-state)
@@
More details regarding the construction of the biased $R_{y}$ gates can be found in [5].
@@
With the possible inclusion of item $m$ we have to update the remaining capacity as well as the total profit – both accounting for the linear profit $p_{m}$ as well as all potentially gained quadratic profits $p_{mm^{\prime}}$. Updating the capacity amounts to subtracting $w_{m}$ from the capacity register’s state, controlled on the $m$-th qubit: $U_{m}^{2}=C_{m}^{1}(\text{SUB}_{w_{m}})$. Analogously, accounting for the linear profit translates into a controlled addition of $p_{m}$ in the profit register $C_{m}^{1}(\text{ADD}_{p_{m}})$. Furthermore – and here lies the main difference between the QTG for ordinary and the Quadratic Knapsack Problem – the quadratic profits have to be taken into an account; for each item $m^{\prime}<m$ such that $p_{mm^{\prime}}>0$ we have to introduce a doubly-controlled addition of $p_{mm^{\prime}}$, respectively: $C_{mm^{\prime}}^{1}(\text{ADD}_{p_{mm^{\prime}}})$. These profit updates are all summarised as
@@
The just described QTG for the QKP now lays the foundation for applying quantum maximum finding (QMaxSearch) [13], tailored to this very problem. More precisely, the QTG enters a state preparation circuit in amplitude amplification (QSearch) [1], which is called as a subroutine in QMaxSearch. Starting with an initial solution and its total profit as initial threshold, we employ an integer comparison oracle on the profit register in order to mark – and subsequently amplify – those states possessing a total profit larger than the current threshold. The threshold is updated throughout the routine with the intermediate solutions’ total profits.
@@
To benchmark the QTG-based search against its best classical competitors, we further expand the high-level simulator introduced by Wilkening et al. [5]. We chose the sampling-based method to assist our simulator. By following the resource estimation techniques from [5], we calculate the number of necessary quantum gates and cycles as well as the success probabilities of obtaining optimal solutions for various MDKP instances.
@@
Figure 2: QTG-based search vs Gurobi: performance analysis. We benchmark the algorithms on relevant benchmark instances with up to $2000$ variables for the QKP and $1500$ variables and $100$ constraints for the MDKP. Each data point represents the time required for Gurobi to obtain an incumbent solution and for the QTG-based search to find an equally good or better solution. The markers’ sizes and colors denote the instance’s size and the solution’s quality, respectively. Assuming a quantum cycle time limit of $1ns$ [14], we observe that for the given benchmark sets the QTG search has the potential to outperform the state-of-the-art solver Gurobi. This is especially noticeable when dealing with problems that are more difficult to solve for classical algorithms (QKP).
@@ IV Results
We demonstrate the capabilities of QTG on a diverse set of benchmark instances. All experiments were carried out on a regular desktop workstation with an AMD Ryzen 7 5800X ($8\times 3.8$ GHz) CPU and $128$ GB of RAM. We slightly alter the benchmarking objective compared to [5]. We compare the solution finding capabilities of QTG with the Gurobi [6] solver. To compare the classical and quantum sides, we use benevolent assumptions for the quantum algorithm explicitly stated in Subsection I-A. To obtain a meaningful runtime for QTG, we assume a quantum cycle time of $1ns$ [14].
@@ Acknowledgment
This work was supported by the Quantum Valley Lower Saxony, the BMBF project QuBRA, and the BMWi project ProvideQ.

### v2

# A quantum search method for quadratic and multidimensional knapsack problems

@@ I Introduction
In recent years, the potential of real-world quantum computing has fueled optimism about solving complex combinatorial optimization problems, prompting extensive theoretical research into developing and analyzing effective quantum methods. So far, the development of new quantum algorithms has largely been driven by theoretical measures, especially worst-case asymptotic runtime analysis [1, 2, 3]. However, a major challenge remains in demonstrating quantum advantage for practical applications. Recent work targeted combinatorial optimization problems that provide a rich set of application areas [4, 5], but also have strong classical competitors: For these problems, classical solvers like Gurobi [6], CPLEX [7] or CP-SAT [8] routinely solve problems involving thousands of variables to provable optimality. An emerging benchmarking discipline [9, 10, 5, 11] is to combine classical emulation techniques, quantitative complexity bounds, and sophisticated, problem- and algorithm-specific simulations to benchmark quantum algorithm against the classical powerhouses for problem instances far larger than the limit of roughly 50 qubits imposed by state vector-based simulation [12].
@@ I-A Assumptions on the Quantum Processing Unit
The resource estimation setup for applying the QTG method to QKP and MDKP relies on several fault-tolerant assumptions for the quantum processing unit (QPU). It assumes that all qubits and gates are noiseless, logical components. The universal gate set includes all single-qubit gates, singly-controlled single-qubit rotations, and Toffoli gates, all of which have equal implementation cost. Disjoint gates can be executed in parallel, counting as a single cycle on the QPU. Singly-controlled single-qubit rotations and Toffoli gates can be applied to arbitrary tuples of qubits. Additionally, all multi-controlled gates have access to the same ancilla qubits and implementing a control on 00 is just as costly as on $1$. Finally, to calculate a runtime from the cycle counts, we assume a quantum cycle time limit of $1ns$ given by [14].
@@ I-B Related Work
The QKP was first introduced by Gallo et al. [15] and has a straightforward quadratic programming formulation (see Section II). Its practical solvability often depends on instance properties such as density [16, 17]. This has drawn significant interest from the optimization community, leading to recent exact algorithms like those in [18, 19] as well as heuristic approaches such as [20, 21]. Other approaches include a quantum-inspired evolutionary algorithms [22]. In contrast, the MDKP tends to be more tractable for integer programming solvers such as Gurobi [23]. Recent approaches, such as those in [24], have proven particularly effective for solving large instances [17]. As for the QKP, heuristic approaches [25, 26, 27] as well as quantum-inspired heuristics [28] have been studied.
Apart from QTG, many other quantum algorithms and frameworks for solving combinatorial optimization problems have been proposed and studied, such as Quantum Annealing [29], [30], [31], [32], Nested Quantum Search (NQS) [33], Quantum Approximate Optimization Algorithm (QAOA) [34] and it’s generalization to the Quantum Alternating Operator Ansatz [35], Quantum Backtracking [36], [37], and Quantum Branch-and-Bound (QBnB) [3]. All of these frameworks are, in principle, also applicable to a variety of knapsack problems. For example, while NQS and QBnB theoretically promise an advantage over classical solvers, Wilkening et al. [5] showed they are significantly outperformed by both the QTG search method and classical solvers on large real-world benchmarks when applied to the KP.
Several variants of QAOA have been to the ordinary KP such as in [38] (introduces a hyperparameter-free objective function and special copula mixers), [39] (combines the QTG with the Grover-mixer QAOA framework [40]), and [41] (combines QAOA with quantum phase estimation techniques [42]). Recently, the scope of QAOA has also been enlarged to the QKP [43] and the MDKP [44].
Another established quantum approach to the QKP is a quantum walk-based genetic algorithm [45] while [46] applied the framework of quantum maximum finding [13] to the MDKP. The latter therefore utilizes the same framework as our approach, but does not incorporate a customized problem-specific state preparation and diffusion circuit like the QTG.
@@ Methods
The QTG $\mathcal{G}$ for the QKP follows the same design criteria as its counterpart for the ordinary 00-$1$ Knapsack Problem [5]. We give here an independent introduction of the QTG, but draw comparisons to it’s original formulation whenever suitable. The entire state preparation part contains $n$ layer unitaries $U_{m}=U_{m}^{3}U_{m}^{2}U_{m}^{1}$ which correspond to including/excluding each item in superposition while keeping track of the remaining capacity and the total profit arising from both the diagonal profits $p_{m}$ and the off-diagonal profits $p_{mm^{\prime}}$. Throughout this process, feasibility is ensured by controlling the branching at each item $m$ on the comparison “$w_{m}\leq c_{\bm{x}}$”. That is, we create a superposition of item $m$ being excluded/included if and only if its weight does not surpass the remaining capacity stored in the second register $\mathcal{H}_{2}$. Correspondingly, if the latter expression is not valid the $m$-th qubit stays in the |$\ket{0}$-state, indicating the item’s exclusion. The corresponding unitary is denoted $U_{m}^{1}=C_{\geq w_{m}}^{2}(R_{y,b}^{x_{m}})$, where the $R_{y}$ gate is responsible for creating a biased superposition of item exclusion (|$\ket{0}$-state) and inclusion (|$\ket{1}$-state)
@@
For $b=0$, we recover the Hadamard gate, mapping |$\ket{0}$ to a uniform superposition of |$\ket{0}$ and |$\ket{1}$. This corresponds to including or excluding the items with equal probability. A non-zero bias $b$ helps maximizing the sampling probability of paths $\bm{x}$ with specific Hamming distance $\Delta\coloneqq\Delta(\bm{x}^{\prime},\bm{x})$ to a given incumbent $\bm{x}^{\prime}$. In [5] it is shown that the optimal choice for $b$ for fixed $\Delta$ is $b_{\text{opt}}\approx n/\Delta$. Hence, for $b\to\infty$, we will recover the incumbent with certainty (as $\Delta\to 0$) while finite values allow us to target specific neighborhoods. Our numerical results indicate that fine-tuning the bias to target distance-4 solutions significantly increases the probability of sampling improved states via the QTG.
@@
With the possible inclusion of item $m$ we have to update the remaining capacity as well as the total profit – both accounting for the linear profit $p_{m}$ as well as all potentially gained quadratic profits $p_{mm^{\prime}}$. Updating the capacity amounts to subtracting $w_{m}$ from the capacity register’s state, controlled on the $m$-th qubit: $U_{m}^{2}=C_{m}^{1}(\text{SUB}_{w_{m}})$, realized e.g. via QFT-based addition circuits [47]. Analogously, accounting for the linear profit translates into a controlled addition of $p_{m}$ in the profit register $C_{m}^{1}(\text{ADD}_{p_{m}})$. Furthermore – and here lies the main difference between the QTG for ordinary and the Quadratic Knapsack Problem – the quadratic profits have to be taken into an account; for each item $m^{\prime}<m$ such that $p_{mm^{\prime}}>0$ we have to introduce a doubly-controlled addition of $p_{mm^{\prime}}$, respectively: $C_{mm^{\prime}}^{1}(\text{ADD}_{p_{mm^{\prime}}})$. These profit updates are all summarised as
@@
The just described QTG for the QKP now lays the foundation for applying quantum maximum finding (QMaxSearch) [13], tailored to this very problem. In a nutshell, QMaxSearch consists of the consecutive application of amplitude amplification (QSearch) [1] where the oracle is given by coherent integer comparison of each solutions’ total profit to the total profit of the incumbent such that solutions with higher profit are amplified. The initial incumbent is obtained via a classical heuristic, e.g. a greedy packing method, and updated whenever the measurement after one round of amplitude amplification yields a better solution. In our case, the QTG enters the amplitude amplification as state preparation circuit as well as part of the diffusion operator.
@@
Figure 2: QTG-based search vs Gurobi: performance analysis. We benchmark the algorithms on relevant benchmark instances with up to $2000$ variables for the QKP and $1500$ variables and $100$ constraints for the MDKP. Each data point represents the time required for Gurobi to obtain an incumbent solution and for the QTG-based search to find an equally good or better solution under benevolent assumptions for the quantum algorithm, see Subsection I-A. The markers’ sizes and colors denote the instance’s size and the solution’s quality, respectively. Assuming a quantum cycle time limit of $1ns$ [14], we observe that for the given benchmark sets the QTG search has the potential to outperform the state-of-the-art solver Gurobi. This is especially noticeable when dealing with problems that are more difficult to solve for classical algorithms (QKP).
@@ IV Results
We demonstrate the capabilities of QTG on a diverse set of benchmark instances. All experiments were carried out on a regular desktop workstation with an AMD Ryzen 7 5800X ($8\times 3.8$ GHz) CPU and $128$ GB of RAM. We slightly alter the benchmarking objective compared to [5]. We compare the solution finding capabilities of QTG with the Gurobi [6] solver. To compare the classical and quantum sides, we use benevolent assumptions for the quantum algorithm explicitly stated in Subsection I-A. We further expand the high-level simulator introduced by Wilkening et al. [5]. We chose the sampling-based method to assist our simulator. By following the resource estimation techniques from [5], we calculate the number of necessary quantum gates and cycles as well as the success probabilities of obtaining optimal solutions. To obtain a meaningful runtime for QTG, we assume a quantum cycle time of $1ns$ [14].
@@ Acknowledgment
This work was supported by the DFG through SFB 1227 (DQ-mat) and Germany’s Excellence Strategy via EXC-2123 QuantumFrontiers 390837967 and PhoenixD, as well as the Quantum Valley Lower Saxony, the BMBF project ATIQ, the BMBF project QuBRA, the BMWi project ProvideQ, and the ERC and the DFG via the project ResourceQ.